OpenStudy (anonymous):

Help need step by step instructions. Will give you medal and become a fan (: A circle is inscribed in a square. Write and simplify an expression for the ratio of the area of the square to the area of the circle. For a circle inscribed in a square, the diameter of the circle is equal to the side length of the square.

OpenStudy (anonymous):

what math are you in O.o

OpenStudy (anonymous):

honors algebra II

OpenStudy (anonymous):

oh that would explain why i don't know anything XD sorry i couldn't help

OpenStudy (anonymous):

its fine!

OpenStudy (anonymous):

Area of square is diameter squared or twice the radius squared: (2r)^2 = 4r^2 Area of the circle is pi times radius squared: (pi)r^2 The ratio of the area of the square to the area of the circle: 4r^2 / (pi)r^2 = 4/pi

OpenStudy (anonymous):

ok that makes sense

OpenStudy (anonymous):

That last equation looks like:\[\frac{ 4 \times \cancel {r ^{2}} }{ \Pi \times \cancel {r ^{2}} }\]

OpenStudy (anonymous):

My computer is having trouble with the Openstudy MathJax. Did that last equation come out ok?

OpenStudy (anonymous):

@shaylynnwalls1 ? Did it?

OpenStudy (anonymous):

yes it did thank you!